I don't buy this story - P2 only briefly burped around GPStime 1291608000 which is around 8pm local time, which is when I was recovering the system.
Today. Jordan talked to Jon Feicht - apparently there is some kind of valve in the TP2 forepump, which only opens ~15-20 seconds after turning the pump on. So the loud sound I was hearing yesterday was just some transient phenomenon. So today morning at ~9am, we turned on TP2. Once again, PTP2 pressure hovered around 500 torr for about 15-20 seconds. Then it started to drop, although both Jordan and I felt that the time it took for the pressure to drop in the range 5 mtorr - 1 mtorr was unusually long. Jordan suspects some "soft-start" feature of the Turbo Pumps, which maybe spins up the pump in a more controlled way than usual after an event like a power failure. Maybe that explains why the pressure dropped so slowly? One thing is for sure - the TP2 controller displayed "TOO HIGH LOAD" yesterday when I tried the first restart (before migrating everything to the older UPS unit). This is what led me to interpret the loud sound on startup of TP2 to indicate some issue with the forepump - as it turns out, this is just the internal valve not being opened.
Anyway, we left TP2 on for a few hours, pumping only on the little volume between it and V4, and PTP2 remained stable at 20 mtorr. So we judged it's okay to open V4. For today, we will leave the system with both TP2 and TP3 backing TP1. Given the lack of any real evidence of a failure from TP2, I have no reason to believe there is elevated risk.
As for prioritising UPS swap - my opinion is that it's better to just replace the batteries in the UPS that has worked for years. We can run a parallel reliability test of the new UPS and once it has demonstrated stability for some reasonable time (>4 months), we can do the swap.
I was able to clear the FAULT indicator on the new UPS by running a "self-test". pressing and holding the "mute" button on the front panel initiates this test according to the manual, and if all is well, it will clear the FAULT indicator, which it did. I'm still not trusting this unit and have left all units powered by the old UPS.
Update 1100 Dec 11: The config remained stable overnight so today I reverted to the nominal config of TP3 pumping the annuli and TP2 backing TP1 which pumps the main volume (through the partially open RV2).
According to the Tripp Lite manual, the FAULT icon indicates "the battery-supported outlets are overloaded." The failure of the TP2 dry pump appears to have caused this. After the dry pump failure, the rising pressure in the TP2 foreline caused TP2's current draw to increase way above its normal operating range. Attachment 1 shows anomalously high TP2 current and foreline pressure in the minutes just before the failure. The critical system-wide failure is that this overloaded the UPS before overloading TP2's internal protection circuitry, which would have shut down the pump, triggering interlocks and auto-notifications.
I've investigated the vacuum controls failure that occurred last night. Here's what I believe happened.
From looking at the system logs, it's clear that there was a sudden loss of power to the control computer (c1vac). Also, the system was actually down for several hours. The syslog shows normal EPICS channel writes (pressure readback updates, etc., and many of them per minute) which suddenly stop at 4:12 pm. There are no error or shutdown messages in the syslog or in the interlock log. The next activity is the normal start-up messaging at 7:39 pm. So this is all consistent with the UPS suddenly failing.
First, there are too many electronics on the 1 kVA UPS. The reason I asked us to buy a dual 208/120V UPS (which we did buy) is to relieve the smaller 120V UPS. I envision moving the turbo pumps, gauge controllers, etc. all to the 5 kVA unit and reserving the smaller 1 kVA unit for the c1vac computer and its peripherals. We now have the dual 208/120V UPS in hand. We should make it a priority to get that installed.
Second, there are 1 Hz "blinker" channels exposed for c1vac and all the slow controls machines, each reporting the machine's alive status. I don't think they're being monitored by any auto-notification program (running on a central machine), but they could be. Maybe there already exists code that could be co-opted for this purpose? There is an MEDM screen displaying the slow machine statuses at Sitemap > CDS > SLOW CONTROLS STATUS, pictured in Attachment 2. This is the only way I know to catch sudden failures of the control computer itself.
I can't find anything in the manual that describes the nature of the FAULT message. In fact, it's not mentioned at all. If the unit detects a fault at its output, I would expect a bit more information. This unit does a programmable level of input error protection, too, usually set at 100%. Still, there is no indication in the manual whether an input issue would be described as a fault; that usually means a short or lifted ground at the output.
Is that a fault code that you can decipher in the manual, or just a light telling you nothing but your UPS is dead?
Around 7pm, the UPS at the vacuum rack seems to have failed. Don't ask me why I decided to check the vacuum screen 10 mins after the failure happened, but the point is, this was a silent failure so the protocols need to be looked into.
Going to the rack, I saw (unsurprisingly) that the 120V UPS was off.
For now, I think this is a safe state to leave the system in. Unless I hear otherwise, I will leave it so - I will be in the lab another hour tonight (~10pm).
Some photos and a screen-cap of the Vac medm screen attached.
As discussed at the meeting, I decided to effect a satellite box swap for the misbehaving MC1 unit. I looked back at the summary pages Vmon for the SRM channels, and found that in the last month or so, there wasn't any significant evidence of glitchiness. So I decided to effect that swap at ~4pm today. The sequence of steps was:
One thing I was reminded of is that the motion of the MC1 optic by controlling the bias sliders is highly cross-coupled in pitch and yaw, it is almost diagonal. If this is true for the fast actuation path too, that's not great. I didn't check it just now.
While I was working on this, I took the opportunity to also check the functionality of the RF path of the IMC WFS. Both WFS heads seem to now respond to angular motion of the IMC mirror - I once again dithered MC2 and looked at the demodulated signals, and see variation at the dither frequency, see Attachment #1. However, the signals seem highly polluted with strong 60 Hz and harmonics, see the zoomed-in time domain trace in Attachment #2. This should be fixed. Also, the WFS loop needs some re-tuning. In the current config, it actually makes the MC2T RIN worse, see Attachment #3 (reference traces are with WFS loop enabled, live traces are with the loop disabled - sorry for the confusing notation, I overwrote the patched version of DTT that I got from Erik that allows the user legend feature, working on getting that back).
The MC1 suspension has begun to show evidence of glitches again, from Friday/Saturday. You can look at the suspension Vmon tab a few days ago and see that the excess fuzz in the Vmon was not there before. The extra motion is also clearly evident on the MCREFL spot. I noticed this on Saturday evening as I was trying to recover the IMC locking, but I thought it might be Millikan so I didn't look into it further. Usually this is symptomatic of some Satellite box issues. I am not going to attempt to debug this anymore.
I suspect what happened here is that the IP didn't get updated when we went from the 131.215.113.xxx system to 192.168.113.xxx system. I fixed it now and can access the web interface. This system is now ready for remote debugging (from inside the martian network obviously). The IP is 192.168.113.90.
Managed to pull this operation off without crashing the RFM network, phew.
BTW, a windows laptop that used to be in the VEA (I last remember it being on the table near MC2 which was cleared sometime to hold the spare suspensions) is missing. Anyone know where this is ?
Continuting the IFO recovery - I am unable to recover similar levels of TRX RIN as I had before. Attachment #1 shows that the TRX RIN is ~4x higher in RMS than TRY RIN (the latter is commensurate with what we had previously). The excess is dominated by some low frequency (~1 Hz) fluctuations. The coherence structure is confusing - why is TRY RIN coherent with IMC transmission at ~2 Hz but not TRX? But anyways, doesn't look like its intensity fluctuations on the incident light (unsurprisingly, since the TRY RIN was okay). I thought it may be because of insufficient low-frequency loop gain - but the loop shape is the same for TRX and TRY. I confirmed that the loop UGF is similar now (red trace in Attachment #2) as it was ~1 month ago (black trace in Attachment #2). Seismometers don't suggest excess motion at 1 Hz. I don't think the modulation depth at 11 MHz is to blame either. As I showed earlier, the spectrum of the error point is comparable now as it was previously.
What am I missing?
The ITMX Oplev (installed in March 2019) was near end of life judging by the SUM channel (see Attachment #1). I replaced it yesterday evening with a new HeNe head. Output power was ~3.25 mW. The head was labelled appropriately and the Oplev spot was recentered on its QPD. The lifetime of ~20 months is short but recall that this HeNe had already been employed as a fiber illuminator at EX and so maybe this is okay.
Loop UGFs and stability margins seem acceptable to me, see Attachment #2-#3.
I updated the ndscope on rossa to a bleeding edge version (0.7.9+dev0) which has many of the fixes I've requested in recent times (e.g. direct PDF export, see Attachment #1). As usual if you find issue, report it on the issue tracker. The basic functionality for looking at signals seems to be okay so this shouldn't adversely impact locking efforts.
In hindsight - I decided to roll-back to 0.7.9, and have the bleeding edge as a separate binary. So if you call ndscope from the command line, you should still get 0.7.9 and not the bleeding edge.
I measured the modulation depth at 11 MHz andf 55 MHz using an optical beat + PLL setup. Both numbers are ~0.2 rad, which is consistent with previous numbers. More careful analysis forthcoming, but I think this supports my claim that the optical gain for the PDH locking loops should not have decreased.
In favor of keeping the same servo gains, I tuned the digital demod phases for the POX and POY photodiode signals to put as much of the PDH error signal in the _I quadrature as possible. The changes are summarized below:
The old locking settings seem to work fine again. This setting isn't set by the ifoconfigure scripts when they do the burt restore - do we want it to be?
Attachments #1 and #2 show some spectra and TFs for the POX/POY loops. In Attachment #2, the reference traces are from the past, while the live traces are from today. In fact, to have the same UGF as the reference traces (from ~1 year ago), I had to also raise the digital servo loop gain by ~20%. Not sure if this can be put down to a lower modulation depth - at least, at the output on the freq ref box, I measured the same output power (at the 0dB variable attenuator gain setting we nominally run in) before and after the changes. But I haven't done an optical measurement of the modulation depth yet. There is also a hint of lesser phase available at ~100 Hz now compared to a year ago.
There seems to be significant phase loss in the TTFSS path, which is limiting the IMC OLTF to <100 kHz.
See Attachment #1 and #2. The former shows the phase loss, while the latter is just to confirm that the optical gain of the error point is roughly the same, since I noticed this after working on and replacing the RF frequency distribution unit. Unfortunately there have been many other changes also (e.g. the work that Rana and Koji did at the IMC rack, swapping of backplane controls etc etc - maybe they have an OLTF measurement from the time they were working?) so I don't know which is to blame. Off the top of my head, I don't see how the RF source can change the phase lag of the IMC servo at 100 kHz. The only part of the IMC RF chain that I touched was the short cable inside the unit that routes the output of the Wenzel source to the front panel SMA feedthrough. I confirmed with a power meter that the power level of the 29.5 MHz signal at that point is the same before and after my work.
The time domain demod monitor point signals appear somewhat noisier in todays measurement compared to some old data I had from 2018, but I think this isn't significant. Once the SR785 becomes available, I will measure the error point spectrum as well to confirm. One thing I noticed was that like many of our 1U/2U chassis units, the feedthrough returns are shorted to the chassis on the RF source box (and hence presumably also to the rack). The design doc for this box makes many statements about the precautions taken to avoid this, but stops short of saying if the desired behavior was realized, and I can't find anything about it in the elog. Can someone confirm that the shields of all the connectors on the box were ever properly isolated? My suspicion is that the shorting is happening where the all-metal N-feedthroughs touch the drilled surfaces on the front panel - while the front and back surfaces of the panel are insulating, the machined surfaces are not.
This is an unacceptable state but no clear ideas of how to troubleshoot quickly (without going piece by piece into the IMC servo chain) occur to me. I still don't understand how the freq source work could have resulted in this problem but I'm probably overlooking something basic. I'm also wondering why the differential receiving at the TTFSS error point did not require a gain adjustment of the IMC servo? Shouldn't the differential-receiving-single-ended-sending have resulted in an overall x0.5 gain?
Update 8 Dec 1200: To test the hypothesis, I bypassed the SR560 based differential receiving and restored the original config. I am then able to run with the original gain settings, and you see in Attachment #4 that the IMC OLTF UGF is back above 100 kHz. It is still a little lower than it was in June 2019, not sure why. There must be some saturation issues somewhere in the signal chain because I cannot preserve the differential receiving and retain 100 kHz UGF, either by raising the "VCO gain" on the MC servo board, setting the SR560 to G=2, or raising the "Common Gain Adjust" on the FSS box by 6 dB. I don't have a good explanation for why this worked for some weeks and failed now - maybe some issue with the SR560? We don't have many working units so I didn't try switching it.
So either there is a whole mess of lines or the frequency noise suppression is limited. Sigh.
This work is now complete. The box was characterized and re-installed in 1X2. I am able to (briefly) lock the IMC and see PDH fringes in POX and POY so the lowest order checks pass.
Even though I did not deliberately change anything in the 29.5 MHz path, and I confirmed that the level at the output is the expected 13 dBm, I had to lower then IN1 gain of the IMC servo by 2dB to have a stable lock - should confirm if this is indeed due to higher optical gain at the IMC error point, or some electrical funkiness. I'm not delving into a detailed loop characterization today - but since my work involved all elements in the RF modulation chain, some detailed characterization of all the locking loops should be done - I will do this in the coming week.
After tweaking the servo gains for the POX/POY loops, I am able to realize the single arm locks as well (though I haven't dont the characterization of the loops yet).
I'm leaving the PSL shutter open, and allowing the IMC autolocker to run. The WFS loops remain disabled for now until I have a chance to check the RF path as well.
Unrelated to this work: Koji's swapping back of the backplane cards seems to have fixed the WFS2 issue - I now see the expected DC readbacks. I didn't check the RF readbacks tonight.
Update 7 Dec 2020 1 pm: A ZHL-2 with heat sink attached and a 11.06 MHz Wenzel source were removed from the box as part of this work (the former was no longer required and the latter wasn't being used at all). They have been stored in the RF electronics cabinet along the east arm.
This turned out to be a much more involved project than I expected. The layout is complete now, but I found several potentially damaged sections of cabling (the stiff cables don't have proper strain relief near the connectors). I will make fresh cables tomorrow before re-installing the unit in the rack. Several changes have been made to the layout so I will post more complete details after characterization and testing.
I was poring over minicircuits datasheets today, and I learned that the minicircuits bandpass filters (SBP10.7 and SBP60) are not bi-directional! The datasheet clearly indicates that the Male SMA connector is the input and the Female SMA connector is the output. Almost all the filters were installed the other way around 😱 . I'll install them the right way around now.
I have the setup built for the AA/AI board testing around the PD testing area. Please let me leave it like that for a week or so.
12/4 TF Tested 5 PCBs
12/6 TF Tested 19 PCBs (12min/PCB) - found 1 failure (S2001479 CH1) -> Fixed 12/11
12/8 TF Tested 16 PCBs (12min/PCB)
PSD Tested 4 PCBs (11min/PCB)
12/11 TF Tested 10 PCBs + 1 fixed channel (All channels checked)
PSD Tested 10 PCBs (11min/PCB)
12/14 PSD Tested 4 PCBs (6.5min/PCB) fixed noise issue of 2 ch, TF issue of 1 ch
12/15 PSD Tested 32 PCBs (6.5min/PCB) fixed noise issue of 1ch
Temp dependence measurement
I checked the backplane connection for IMC WFS2 and found that the cables for IMC WFS2 and the IMC demod were swapped during my IMC noise hunting activities. I reverted it just now.
But we need to check if this damaged anything such as the WFS2 head, the WFS2 demod, etc, once the IMC locking is back.
1. That's true. But we are already in that regime with the Var attn at 0dB, aren't we? We can reduce the input to the amp by 1-2dBm sacrificing the EOM out by that amount (we can compensate this for the demo out by removing the 1dB attn).
2. Not 100% sure but one possible explanation is that we wanted to keep the Marconi output large (or as large as possible) to keep the SNR between the signal and the noise of the driver in Marconi. The attenuator is less noisy compared to the driver noise.
3. My guess is that theoretically we were supposed to have 13dBm input and 20dBm output in design. However, the actual input was as such. We can restore it to the 13dBm input.
I'm open to either approach. If the full replacement requires a lot of machining, maybe I will stick to just the 55 MHz line. But if only a couple of new holes are required, it might be advantageous to do the revamp while we have the box out? What do you think?
BTW, now that I look more closely at the RF chain, I have several questions:
I guess it is feasible to have +17 dBm of 55 MHz signal to plug into the Quad Demod chassis - e.g. drive the 55 MHz input with 20 dBm, pick off 3dBm to the front panel for ASC. Then we can even have several "spare" 55 MHz outputs and still satisfy the 9 dBm input that the ZHL-2 in the 55 MHz chain wants (though again, isn't this dangerously close to the 1dB compression point?). The design doc claims to have done some Optickle modeling, so I guess there isn't really any issue?
Are you going to full replacement of the 55MHz system? Or just remove the 7dBm and then implement the proposed modification for the 55MHz line?
Let's use RG405 for better shielding. It is not too stiff. The bending (just once) does not break the cable.
I removed the Frequency Generation box from the 1X2 rack. For the time being, the PSL shutter is closed, since none of the cavities can be locked without the RF modulation source anyways.
Prior to removal, I did the following:
One thing I noticed was that we're using very stiff coax cabling (RG405) inside this box? Do we need to stick with this option? Or can we use the more flexible RG316? I guess RG405 is lower loss, so it's better. I can't actually find any measurement of the shielding performance in my quick google searching but I think the claim on the call yesterday was that RG405 with its solder soaked braids offer superior shielding.
Before doing any modification you should check how much the distributed powers are at the ports.
Also your modification will change the relative phase between 11MHz and 55MHz.
Can you characterize how much phase difference you have between them, maybe using the modulation of the main marconi? And you might want to adjust it to keep the previous value (or any new value) after the modification by adding a cable inside?
Update to the gauge replacement plan (15692), based on Jordan's walk-through today. He confirmed:
Based on this info (and also info from Gautam that the PAN gauge is still working), I've updated the plan as follows. In summary, I now propose we install the fifth FRG in the TP1 foreline (PTP1 location) and leave P2 and P3 where they are, as they are no longer needed elsewhere. Any comments on this plan? I plan to order all the necessary gaskets, blanks, etc. tomorrow.
I got a bit confused by your description.
The demod board claims that the nominal power at each LO port is 10dBm. So we want to give at least 16dBm to the (external?) 4way power splitter, but we only have 15dBm. As you said, the actual LO power reaching the FET mixier (PE4140) is the level of ~20dBm. But you said the requirement for the mixer is -7dBm. So are you proposing to reduce the LO level (slightly) than the LIGO recommendation because the minimum for PE4140 is -7dBm?
If that's the message, then I can say "yes". We supply 8~9dBm to the LO ports instead of 10dBm. I suppose the mixers don't care about this level of reduction.
Looking at my original post [40m ELOG 11817], the necessary modification is much larger than you have indicated in your post (as yours is the modification of my modification plan.)
If you do your modification you have to deal with the components rearrangement in the chassis. I think you can still accomplish it as you are going to remove an amplifier and gain the space from it.
The main RF line still has 5dBm Attn. How about to insert another 3dB power splitter there and create a spare 55MHz port for the future use?
Before doing any modification you should check how much the distributed powers are at the ports.
Also your modification will change the relative phase between 11MHz and 55MHz.
Can you characterize how much phase difference you have between them, maybe using the modulation of the main marconi? And you might want to adjust it to keep the previous value (or any new value) after the modification by adding a cable inside?
looks good to me.
The thing I usually look for is how much the downstream system (mixers, etc) can perturb the main oscillator. i.e. we don't want mixer in one chain to reflect back and disturb the EOM chain. But since our demods have amplifiers on the LO side we're pretty immune to that.
Since we will have several new 1U / 2U aLIGO style electronics chassis installed in the racks, it is desirable to have a more compact power distribution solution than the fusable terminal blocks we use currently.
I did a quick walkaround of the lab and the electronics rack today. I estimate that we will need 5 units of the 24 V and 5 units of the 18 V power strips. Each end will need 1 each of 18 V and 24 V strips. The 1Y1/1Y2/1Y3 (LSC/OMC/BHD sus) area will be served by 1 each 18 V and 24 V. The 1X1/1X2 (IOO) area will be served by 1 each 18 V and 24 V. The 1X5/1X6 (SUS Shadow sensor / Coil driver) area will be served by 1 each of 18 V and 24 V. So I think we should get 7 pcs of each to have 2 spares.
Most of the chassis which will be installed in large numbers (AA, AI, whitening) supports 24V DC input. A few units, like the WFS interface head, OMC driver, OMC QPD interface, require 18V. It is not so clear what the input voltage for the Satellite box and Coil Drivers should be. For the former, an unregulated tap-off of the supply voltage is used to power the LT1021 reference and a transistor that is used to generate the LED drive current for the OSEMs. For the latter, the OPA544 high current opamp used to drive the coil current has its supply rails powered by again, an unregulated tap-off of the supply voltage. Doesn't seem like a great idea to drive any ICs with the unregulated switching supply voltage from a noise point of view, particularly given the recent experience with the HV coil driver testing and the PSRR, but I think it's a bit late in the game to do anything about this. The datasheet specs ~50 dB of PSRR on the negative rail, but we have a couple of decoupling caps close to the IC and this IC is itself in a feedback loop with the low noise AD8671 IC so maybe this won't be much of an issue.
For the purposes of this discussion, I think both Satellite Amp and Coil Driver chassis can be driven with +/- 24 V DC.
On a side note - after the upgrade will the "Satellite Amplifiers" be in the racks, and not close to the flange as they currently are? Or are we gonna have some mini racks next to the chambers? Not sure what the config is at the sites, and if the circuits are designed to drive long cables.
The latest greatest UPS has been delivered. I will move it to near the vacuum rack in its packaging for storage. It weighs >100lbs so care will have to be taken when installing - can the rack even support this?
I'm thinking of making some modifications to the RF distribution box in 1X2, so as to have an extra 55 MHz pickoff. Koji already proposed some improvements to the layout in 2015. I've marked up his "Possible Improvement" page of the document in Attachment #1, with my proposed modifications. I believe it will be possible to get 15-16 dBm of signal into a 4 way RF splitter in the quad demod chassis. With the insertion loss of the splitter, we can have 9-10 dBm of LO reaching each demod board, which will then be boosted to +20 dBm by the Teledyne on board. The PE4140 mixer claims to require only -7 dBm of LO signal. So we have quite a bit of headroom here - as long as we limit the RF signal to 0dBm (=0.5 Vpp from the LMH6431 opamp at 55 MHz, we shouldn't be having a much larger signal anyways), we should be just fine with 15 dBm of LO power (which is what we will have after the division into the I and Q paths, and nominal insertion losses in the transmission path). These numbers may be slight overestimates given the possible degradation of the RF amps over the last 10 years, but shouldn't be a show-stopper.
Do the RF electronics experts agree with my assessment? If so, I will start working on these mods tomorrow. Technically, the splitter can be added outside the box, but it may be neater if we package it inside the box.
The summary pages were in a sad state of disrepair - the daily jobs haven't been running for > 1 month. I only noticed today because Jordan wanted to look at some vacuum trends and I thought summary pages is nice for long term lookback. I rebooted it just now, seems to be running. @Tega, maybe you want to set up some kind of scripted health check that also sends an alert.
As discussed at the meeting, I commenced the recovery of the CDS status at 1750 local time.
Single arm POX/POY locking was checked, but not much more. Our IMC WFS are still out of service so I hand aligned the IMC a bit, IMC REFL DC went from ~0.3 to ~0.12, which is the usual nominal level.
FYI, there is this. Seems pretty well maintained, and so might be more useful in the long run. The available catalog of instruments is quite impressive - TC200 temp controller and SRS345 func gen are included and are things we use in the lab. maybe you can make a pull request to add MDT694B (there is some nice API already built I think). We should also put our netgpibdata stuff and the vacuum gauge control (basically everything that isn't rtcds) on there (unless there is some intellectual property rights issues that the Caltech lawyers have to sort out).
Given the similarities between the MDT694B (single channel piezo controller) and TC200 (temperature controller) serial interfaces, I added the pyserial driver here.
*Warning* this first version of the driver remains untested
Now that the new Agilent full-range gauges (FRGs) have been received, I'm putting together an installation plan. Since my last planning note in Sept. (ELOG 15577), two more gauges appear to be malfunctioning: CC2 and PAN. Those are taken into account, as well. Below are the proposed changes for all the sensors in the system.
For replacements, I recommend we consider the Agilent FRG-700 Pirani Inverted Magnetron Gauge. It uses dual sensing techniques to cover a broad pressure range from 3e-9 torr to atmosphere in a single unit. Although these are more expensive, I think we would net save money by not having to purchase two separate gauges (Pirani + hot/cold cathode) for each location. It would also simplify the digital controls and interlocking to have a streamlined set of pressure readbacks.
For controllers, there are two options with either serial RS232/485 or Ethernet outputs. We probably want the Agilent XGS-600, as it can handle all the gauges in our system (up to 12) in a single controller and no new software development is needed to interface it with the slow controls.
I don't think your simulation looked inaccurate (at least not to me). In my opinion, we just want to minimize any excess noise from intermodulation. Of course, its possible that stuffing too many notches will make it difficult to have the same low noise as a simple circuit, so that's worth considering.
Also, the intermodulation is mainly a problem when the other peaks are not suppressed by some feedback: e.g. POP55_I can have excess noise if POP55_Q or POP11_I are not controlled by some MICH/PRCL/SRCL loops.
For the WFS, perhaps this is not a significant issue, but I'm not sure. My suggestion is to stuff 11 & 55 for sure, and then the others depending on the amplitude of the peaks and the consequent intermodulation. IF it works with all stuffed, that seems good. If its tricky to get it to work with all stuffed, I'd back off on a couple of them...but it probably takes more careful thought to figure out which ones are least important.
An 8 channel whitening chassis was prepared and tested. I measured:
Whitening chassis. Waiting for front panels to arrive, PCBs and interface board are in hand, stuffed and ready to go. A question here is how we want to control the whitening - it's going to be rather difficult to have fast switchable whitening. I think we can just fix the whitening state. Another option would be to control the whitening using Acromag BIO channels.
I am confused by the discussion during the call today. I revisited Hartmut's paper - the circuit in Fig 6 is essentially what I am calling "only 2f_2 notch stuffed" in my previous elog. Qualitatively, the plot I presented in Attachment #2 of the preceeding elog in this thread shows the expected behavior as in Fig 8 of the paper - the impedance seen by the photodiode is indeed lower. In Attachment #1, I show the comparison - the "V(anode)/I(I1)" curve is analogous to the "PD anode" curve in Hartmut's paper, and the "V(vout)/I(I1)" curve is analogous to the "1f-out" curve. I also plot the sensitivity analysis (Attachment #2), by varying the photodiode junction capacitance between 100pF and 200 pF (both values inclusive) in 20 pF steps. There is some variation at 55 MHz, but it is unlikely that the capacitance will change so much during normal operation?
I understand the motivation behind stuffing the other notches, to reduce intermodulation effects. But the impression I got from the call was that somehow, the model I presented was wrong. Can someone help me identify the mistake?
I didn't bother to export the LTspice data and make a matplotlib plot for this quick analysis, so pardon the poor presentation. The colors run from green=100pF to grey=200pF.
On the call last week, I claimed that there isn't much hope of directly measuring Ponderomotive Squeezing in aLIGO without some significant configurational changes. Here, I attempt to quantify this statement a bit, and explicitly state what I mean by "significant configurational changes".
The I/O relations will generally look something like:
The. magnitudes of the matrix elements C_12 and C_21 (i.e. phase to amplitude and amplitude to phase coupling coefficients) will encode the strength of the Ponderomotive squeezing.
For the inital study, let's assume DC readout (since there isn't a homodyne readout yet even in Advanced LIGO). This amounts to setting in the I/O relations, where the former angle is the "homodyne phase" and the latter is the "SRC detuning". For DC readout, the LO quadrature is fixed relative to the signal - for example, in the usual RSE operation, . So the quadrature we will read out will be purely (or nearly so, for small detunings around RSE operation). The displacement noises will couple in via the matrix element. Attachment #1 and Attachment #2 show the off-diagonal elements of the "C" matrix for detunings of the SRC near RSE and SR operation respectively. You can see that the optomechanical coupling decays pretty rapidly above ~40 Hz.
In this particular case, there is no benefit to detuning the SRC, because we are assuming the homodyne angle is fixed, which is not an unreasonable assumption as the quadrature of the LO light is fixed relative to the signal in DC readout (not sure what the residual fluctuation in this quantity is). But presumably it is at the mrad level, so the pollution due to the orthogonal anti-squeezed quadrture can be ignored for a first pass I think. I also assume ~10 degrees of detuning is possible with the Finesse ~15 SRC, as the linewidth is ~12 degrees.
To see how this would look in an actual measurement, I took the data from Lee's ponderomotive squeezing paper, as an estimate for the classical noises, and plotted the quantum noise models for a few representative SRC detunings near RSE operation - see Attachment #3. The curves labelled for various phis are the quantum noise models for those SRC detunings, assuming DC readout. I fudged the power into the IFO to make my modelled quantum noise curve at RSE line up with the high frequency part of the "Measured DARM" curve. To measure Ponderomotive Squeezing unambiguously, we need the quantum noise curve to "dip" as is seen around 40 Hz for an SRC tuning of 80 degrees, and that to be the dominant noise source. Evidently, this is not the case.
The case for balanced homodyne readout:
I haven't analyzed it in detail yet - but it may be possible that if we can access other quadratures, we might benefit from rotating away from the DARM quadrature - the strength of the optomechanical coupling would decrease, as demonstrated in Attachments #1 and #2, but the coupling of classical noise would be reduced as well, so we may be able to win overall. I'll briefly investigate whether a robust measurement can be made at the site once the BHD is implemented.
Last week and this week I've been working on the characterization of the Q3000 QPDs. The QPDs were named 81, 82, 83, and 94.
My recommendation is to use #81 and #84 as they have similar dark current characteristics between the segments. But basically, all the QPDs look fine.
The actual junction capacitance and the RF dark noise should be characterized by the actual WFS head circuit.
The QPD packages were labeled and returned to Gautam to be implemented in the WFS heads.
gautam: S/N #84 was installed as the AS WFS QPD. The remaining 3 are stored in the clean cabinet at EX (where the rest of the RF photodiodes are).
Five Agilent pressure gauges were delivered to the 40m. It is stored with the controller and cables in the office area. This completes the inventory for the gauge replacement - we have all the ordered parts in hand (though. not necessarily all the adaptor flanges etc). I'll see if I can find some cabinet space in the VEA to store these, the clutter is getting out of hand again...
in addition, the spare gate valve from LHO was also delivered today to the 40m. It is stored at EX with the other spare valves.
It is stored along with the cables that arrived a few weeks ago, awaiting the gauges which are now expected next week sometime.
Basically, they repeated our specs and showed the coating performances for HR/AR for 10deg P and PR/AR for 45deg P. There is no RoC measurement by the vendor.
Nevertheless, their RoC (paper) specs should be compared with our request.
Optics --> Cabinet at south end (Attachment #1)
Scanned datasheets--> wiki. It would be good if someone can check the specs against what was ordered.
Attachment #1 - Proposed mods for 40m RF freqs.
Attachment #2 - Modelled TFs for the case where all the notches are stuffed, and where only the 2f notch is stuffed.
Attachment #3 - Modelled TFs for the case where all the notches are stuffed, and where only the 2f notch is stuffed.
Any other red flags anyone sees before I finish stuffing the board?
WFS head and housing. Need to finalize the RF transimpedance gain (i.e. the LC resonant part), and also decide which notches we want to stuff.
Where do we want to install the interface and readout electronics for the AS port WFS? Options are:
There isn't much difference in terms of cable length that will be required - I believe the AS WFS is going to go on the AP table even in the new optical layout and not on the ITMY in-air oplev table?
The project requires a large number of new electronics modules. Here is a short update and some questions I had:
Approximately half of the assembly of the various electronics is now complete. The basic electrical testing of the interface chassis and demod chassis are also done (i.e. they get power, the LEDs light up, and are stable for a few minutes). Detailed noise and TF characterization will have to be done.
Ordered 11/16 from CDW, on PO# S492940, the high voltage Tripp Lite SMART5000XFMRXL for TP-1. Should be arriving in about a week.
yes, both problems can be fixed. Usually we just order some spare lead-acid batteries from SRS (Steve may have some spare ones somewhere). The DC offset often comes from a busted FET input. I bought 50 of those at one point - they're obsolete. Its also possible to replace the input stage with any old FET pair.
I'll handle the one with the offset if you leave it on my desk.
Shruti picked it up @4pm.
true. also try to choose a cap with a goow high frequency response. In the Electronics Noise book by Ott there's some graph about this. I bet you good do a Bing search and also find something more modern. Basically we want to make sure that the self resonance is not happening at low frequencies. Might be tought to find one with a good HF response, a high voltage rating, and > 1uF.
Yes. The datasheet has a recommendation circuit with 10uF caps. Companies are careful to show reproducible, reliably functional circuit examples on datasheets. So, if the caps are there you should try to replicate the design.
Additional bypass capacitors? I use 0.1 uF, 700V DC ceramic capacitors as bypass capacitors close to the leads of the PA95, as is recommended in the datasheet. Can adding a 10uF capacitor in parallel provide better filtering? I'm not sure if one with compatible footprint and voltage rating is readily available, I'll look around.
I now think the excess noise in this circuit could be coming from the KEPCO switching power supply (in fact, the supplies are linear, and specd for a voltage ripple at the level of <0.002% of the output - this is pretty good I think, hard to find much better).
All component references are w.r.t. the schematic. For this test, I decided to stuff a fresh channel on the board, with new components, just to rule out some funky behavior of the channel I had already stuffed. I decoupled the HV amplifier stage and the Acromag DAC noise filtering stages by leaving R3 open. Then, I shorted the non-inverting input of the PA95 (i.e. TP3) to GND, with a jumper cable. Then I measured the noise at TP5, using the AC coupling pomona box (although in principle, there is no need for this as the DC voltage should be zero, but I opted to use it just in case). The characteristic bump in the spectra at ~100Hz-1kHz was still evident, see the bottom row of Attachment #1. The expected voltage noise in this configuration, according to my SPICE model, is ~10 nV/rtHz, see the analysis note.
As a second test, I decided to measure the voltage noise of the power supply - there isn't a convenient monitor point on the circuit to directly probe the +/- HV supply rails (I didn't want any exposed HV conductors on the PCB) - so I measured the voltage noise at the 3-pin connector supplying power to the 2U chassis (i.e. the circuit itself was disconnected for this measurement, I'm measuring the noise of the supply itself). The output is supposedly differential - so I used the SR785 input "Float" mode, and used the Pomona AC coupling box once again to block the large DC voltage and avoid damage to the SR785. The results are summarized in the top row of Attachment #1.
The shape of the spectra suggests to me that the power supply noise is polluting the output noise - Koji suggested measuring the coherence between the channels, I'll try and do this in a safe way but I'm hesitant to use hacky clips for the High Voltage. The PA95 datasheet says nothing about its PSRR, and seems like the Spice model doesn't include it either. It would seem that a PSRR of <60dB at 100 Hz would explain the excess noise seen in the output. Typically, for other Op-Amps, the PSRR falls off as 1/f. The CMRR (which is distinct from the PSRR) is spec'd at 98 dB at DC, and for other OpAmps, I've seen that the CMRR is typically higher than the PSRR. I'm trying to make a case here that it's not unreasonable if the PA95 has a PSRR <= 60dB @100 Hz.
So what are the possible coupling mechanisms and how can we mitigate it?
What do the analog electronics experts think? I may be completely off the rails and imagining things here.
Update 2130: I measured the coherence between the positive supply rail and the output, under the same conditions (i.e. HV stage isolated, input shorted to ground). See Attachment #2 - the coherence does mirror the "bump" seen in the output voltage noise - but the coherence is. only 0.1, even with 100 averages, suggesting the coupling is not directly linear - anyways, I think it's worth it to try adding some extra decoupling, I'm sourcing the HV 10uF capacitors now.